OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation 8b + 12 = 5(4 + 2b) is given below. Choose the step that is missing from the solution. 8b + 12 = 5(4 + 2b) 8b = 8 + 10b −2b = 8 b = −4 8b + 12 = 20 + 2b 8b + 12 = 20 + 10b 8b + 12 = 9 + 7b 8b + 12 = 4 + 10b

OpenStudy (anonymous):

8b + 12 = 5(4+2b) For the next step, you are to use the distributive property for 5(4+2b). The distributive property is a(b+c) = ab+ac. Can you solve that for 5(4+2b)?